OpenStudy (anonymous):

what polynomial has roots -4, 2, 5

OpenStudy (saifoo.khan):

x = -4 x = 2 x = 5 can be written as, x + 4 x - 2 x - 5 Now.

OpenStudy (saifoo.khan):

(x+4)(x-2)(x-5) = 0 Solve.

OpenStudy (anonymous):

\( f(x)=(x+4)(x-2)(x-5)= 40-18 x-3 x^2+x^3\)